### v2

# Yifan Wang1∗ Jianjun Zhou123∗ Haoyi Zhu1 Wenzheng Chang1 Yang Zhou1

Abstract
We introduce $\pi^{3}$, a feed-forward neural network that offers a novel approach to visual geometry reconstruction, breaking the reliance on a conventional fixed reference view. Previous methods often anchor their reconstructions to a designated viewpoint, an inductive bias that can lead to instability and failures if the reference is suboptimal. In contrast, $\pi^{3}$ employs a fully permutation-equivariant architecture to predict affine-invariant camera poses and scale-invariant local point maps without any reference frames. This design not only makes our model inherently robust to input ordering, but also leads to higher accuracy and performance. These advantages enable our simple and bias-free approach to achieve state-of-the-art performance on a wide range of tasks, including camera pose estimation, monocular/video depth estimation, and dense point map reconstruction. Code and models are publicly available.
Figure 1: $\pi^{3}$ effectively reconstructs a diverse set of open-domain images in a feed-forward manner, encompassing various scenes such as indoor, outdoor, and aerial-view, as well as cartoons, with both dynamic and static content.
1 Introduction
Visual geometry reconstruction, a long-standing and fundamental problem in computer vision, holds substantial potential for applications such as augmented reality [7], robotics [50], and autonomous navigation [17]. While traditional methods addressed this challenge using iterative optimization techniques like Bundle Adjustment (BA) [11], the field has recently seen remarkable progress with feed-forward neural networks. End-to-end models like DUSt3R [39] and its successors have demonstrated the power of deep learning for reconstructing geometry from image pairs [13, 46], videos, or multi-view collections [42, 47, 34].
Despite these advances, a critical limitation persists in both classical and modern approaches: the reliance on selecting a single, fixed reference view. The camera coordinate system of this chosen view is treated as the global frame of reference, a practice inherited from traditional Structure-from-Motion (SfM) [11, 4, 24, 20] or Multi-view Stereo (MVS) [9, 25]. We contend that this design choice introduces an unnecessary inductive bias that fundamentally constrains the performance and robustness of feed-forward neural networks. As we demonstrate empirically, this reliance on an arbitrary reference makes existing methods, including the state-of-the-art (SOTA) VGGT [34], highly sensitive to the initial view selection. A poor choice can lead to a dramatic degradation in reconstruction quality, hindering the development of robust systems (Fig. 2).
To overcome this limitation, we introduce $\pi^{3}$, a robust, accurate, and fully permutation-equivariant method that eliminates reference view-based biases in visual geometry learning. $\pi^{3}$ accepts varied inputs—including single images, video sequences, or unordered image sets from static or dynamic scenes—without designating a reference view. Instead, our model predicts an affine-invariant camera pose and a scale-invariant local pointmap, with the pointmap being defined in that frame’s own camera coordinate system. By eschewing order-dependent components like frame index positional embeddings and employing a transformer architecture that alternates between view-wise and global self-attention (similar to [34]), $\pi^{3}$ achieves true permutation equivariance. This guarantees a consistent one-to-one mapping between visual inputs and the reconstructed geometry, making the model inherently robust to input order and immune to the reference view selection problem (Table 8).
Figure 2: Performance comparison across different reference frames. While previous methods, even with DINO-based selection, show inconsistent results, $\pi^{3}$ consistently delivers superior and stable performance, demonstrating its robustness.
Our design yields significant advantages. Primarily, it is substantially more robust. Unlike previous methods, our approach demonstrates minimal performance degradation and a low standard deviation when the reference frame is altered (Fig. 2 and Table 7). Furthermore, it enhances reconstruction accuracy over earlier methods that rely on a reference view.
Through extensive experiments, $\pi^{3}$ establishes a new SOTA across numerous benchmarks and tasks. For example, it achieves comparable performance to existing methods like MoGe [37] in monocular depth estimation, and outperforms VGGT [34] in video depth estimation and camera pose estimation. On the Sintel benchmark, $\pi^{3}$ reduces the camera pose estimation ATE from VGGT’s 0.167 down to 0.074 and improves the scale-aligned video depth absolute relative error from 0.299 to 0.233. Furthermore, $\pi^{3}$ is both lightweight and fast, achieving an inference speed of 57.4 FPS compared to DUSt3R’s 1.25 FPS and VGGT’s 43.2 FPS. Its ability to reconstruct both static and dynamic scenes makes it a robust and optimal solution for real-world applications.
In summary, the contributions of this work are as follows:
  • We are the first to systematically identify and challenge the reliance on a fixed reference view in visual geometry reconstruction, demonstrating how this common design choice introduces a detrimental inductive bias that limits model robustness and performance.
  • We propose $\pi^{3}$, a novel, fully permutation-equivariant architecture that eliminates this bias. Our model predicts affine-invariant camera poses and scale-invariant pointmaps in a purely relative, per-view manner, completely removing the need for a global coordinate system.
  • We demonstrate through extensive experiments that $\pi^{3}$ establishes a new state-of-the-art on a wide range of benchmarks for camera pose estimation, monocular/video depth estimation, and pointmap reconstruction, outperforming prior leading methods.
2 Related Work
2.1 Traditional 3D Reconstruction
Reconstructing 3D scenes from images is a foundational problem in computer vision. Classical methods, such as Structure-from-Motion (SfM) [11, 4, 24, 20] and Multi-View Stereo (MVS) [9, 25], have achieved considerable success. These techniques leverage the principles of multi-view geometry to establish feature correspondences across images, from which they estimate camera poses and generate dense 3D point clouds. Although robust, particularly in controlled environments, these methods typically rely on complex, multi-stage pipelines. Moreover, they often involve time-consuming iterative optimization problems, such as Bundle Adjustment (BA), to jointly refine the 3D structure and camera poses.
2.2 Feed-Forward 3D Reconstruction
Recently, feed-forward models have emerged as a powerful alternative, capable of directly regressing the 3D structure of a scene from a set of images in a single pass. Pioneering efforts in this domain, such as Dust3R [39], focused on processing image pairs to predict a point cloud within the coordinate system of the first camera. While effective for two views, scaling this to larger scenes requires a subsequent global alignment step, a process that can be both time-consuming and prone to instability.
Subsequent work has focused on overcoming this limitation. Fast3R [42] represents a significant advance by enabling simultaneous inference on thousands of images, thereby eliminating the need for a costly and fragile global alignment stage. Other approaches have explored simplifying the learning problem itself. For instance, FLARE [47] decomposes the task by first predicting camera poses and then estimating the scene geometry. VGGT [34] leverages multi-task learning and large-scale datasets to achieve superior accuracy and performance.
A unifying characteristic of these methods is their reliance on anchoring the predicted 3D structure to a designated reference frame. Our work departs from this paradigm by presenting a fundamentally different approach.
3 Method
Figure 3: Unlike prior methods that designate a reference view by concatenating a special token (Type A) or adding a learnable embedding (Type B), $\pi^{3}$ achieves permutation equivariance by eliminating this requirement altogether. Instead, it employs relative supervision, making our approach inherently robust to the order of input views.
3.1 Permutation-Equivariant Architecture
To ensure our model’s output is invariant to the arbitrary ordering of input views, we designed our network $\phi$ to be permutation-equivariant.
Let the input be a sequence of $N$ images, $S=(\mathbf{I}_{1},\dots,\mathbf{I}_{N})$, where each image $\mathbf{I}_{i}\in\mathbb{R}^{H\times W\times 3}$. The network $\phi$ maps this sequence to a corresponding tuple of output sequences:
$\phi(S)=\mathopen{}\mathclose{{\left((\mathbf{T}_{1},\dots,\mathbf{T}_{N}),(\mathbf{X}_{1},\dots,\mathbf{X}_{N}),(\mathbf{C}_{1},\dots,\mathbf{C}_{N})}}\right)$ (1)
Here, $\mathbf{T}_{i}\in SE(3)\subset\mathbb{R}^{4\times 4}$ is the camera pose, $\mathbf{X}_{i}\in\mathbb{R}^{H\times W\times 3}$ is the associated pixel-aligned 3D point map represented in its own camera coordinate system, and $\mathbf{C}_{i}\in\mathbb{R}^{H\times W}$ is the confidence map of $\mathbf{X}_{i}$, each corresponding to the input image $\mathbf{I}_{i}$.
For any permutation $\pi$, let $P_{\pi}$ be an operator that permutes the order of a sequence. The network $\phi$ satisfies the permutation-equivariant property:
$\phi(P_{\pi}(S))=P_{\pi}(\phi(S))$ (2)
This means that permuting the input sequence, $P_{\pi}(S)=(\mathbf{I}_{\pi(1)},\dots,\mathbf{I}_{\pi(N)})$, results in an identically permuted output tuple:
$\begin{split}P_{\pi}(\phi(S))=\big{(}&(\mathbf{T}_{\pi(1)},\dots,\mathbf{T}_{\pi(N)}),\\ &(\mathbf{X}_{\pi(1)},\dots,\mathbf{X}_{\pi(N)}),\\ &(\mathbf{C}_{\pi(1)},\dots,\mathbf{C}_{\pi(N)})\big{)}\end{split}$ (3)
This property guarantees a consistent one-to-one correspondence between each image and its respective output (e.g., geometry or pose). This design offers several key advantages. First, reconstruction quality becomes independent of the reference view selection, in contrast to prior methods that suffer from performance degradation when the reference view changes. Second, the model becomes more robust to uncertain or noisy observations. These claims are empirically validated in Section 4.
To realize this equivariance in practice, our implementation (illustrated in Fig. 3) omits all order-dependent components. Specifically, we discard all order-dependent components, such as positional embeddings used to differentiate between frames and specialized learnable tokens that designate a reference view, like the camera tokens found in VGGT [34]. Our pipeline begins by embedding each view into a sequence of patch tokens using a DINOv2 [18] backbone. These tokens are then processed through a series of alternating view-wise and global self-attention layers, similar to [34], before a final decoder generates the output. The detailed architecture of our model is provided in Appendix A.1.
3.2 Scale-Invariant Local Geometry
For each input image $\mathbf{I}_{i}$, our network predicts the geometry as a pixel-aligned 3D point map $\hat{\mathbf{X}}_{i}$. Each point cloud is initially defined in its own local camera coordinate system. A well-known challenge in monocular reconstruction is the inherent scale ambiguity. To address this, our network predicts the point clouds up to an unknown, yet consistent, scale factor across all $N$ images of a given scene.
Consequently, the training process requires aligning the predicted point maps, $(\hat{\mathbf{X}}_{1},\dots,\hat{\mathbf{X}}_{N})$, with the corresponding ground-truth (GT) set, $(\mathbf{X}_{1},\dots,\mathbf{X}_{N})$. This alignment is accomplished by solving for a single optimal scale factor, $s^{*}$, which minimizes the depth-weighted L1 distance across the entire image sequence. The optimization problem is formulated as:
$s^{*}=\underset{s}{\arg\min}\sum_{i=1}^{N}\sum_{j=1}^{H\times W}\frac{1}{z_{i,j}}\|s\hat{\mathbf{x}}_{i,j}-\mathbf{x}_{i,j}\|_{1}$ (4)
Here, $\hat{\mathbf{x}}_{i,j}\in\mathbb{R}^{3}$ denotes the predicted 3D point at index $j$ of the point map $\hat{\mathbf{X}}_{i}$. Similarly, $\mathbf{x}_{i,j}$ is its ground-truth counterpart in $\mathbf{X}_{i}$. The term $z_{i,j}$ is the ground-truth depth, which is the z-component of $\mathbf{x}_{i,j}$. This problem is solved using the ROE solver proposed by [37].
Finally, the point cloud reconstruction loss, $\mathcal{L}_{\text{points}}$, is defined using the optimal scale factor $s^{*}$:
$\mathcal{L}_{\text{points}}=\frac{1}{3NHW}\sum_{i=1}^{N}\sum_{j=1}^{H\times W}\frac{1}{z_{i,j}}\|s^{*}\hat{\mathbf{x}}_{i,j}-\mathbf{x}_{i,j}\|_{1}$ (5)
To encourage the reconstruction of locally smooth surfaces, we also introduce a normal loss, $\mathcal{L}_{\text{normal}}$. For each point in the predicted point map $\hat{\mathbf{X}}_{i}$, its normal vector $\hat{\mathbf{n}}_{i,j}$ is computed from the cross product of the vectors to its adjacent neighbors on the image grid. We then supervise these normals by minimizing the angle between them and their ground-truth counterparts $\mathbf{n}_{i,j}$:
$\mathcal{L}_{\text{normal}}=\sum_{i=1}^{N}\sum_{j=1}^{H\times W}\arccos\mathopen{}\mathclose{{\left(\hat{\mathbf{n}}_{i,j}\cdot\mathbf{n}_{i,j}}}\right)$ (6)
We supervise the predicted confidence map $\mathbf{C}_{i}$ using a Binary Cross-Entropy (BCE) loss, denoted $\mathcal{L}_{\text{conf}}$. The ground-truth target for each point is set to 1 if its L1 reconstruction error, $\frac{1}{z_{i,j}}\|s^{*}\hat{\mathbf{x}}_{i,j}-\mathbf{x}_{i,j}\|_{1}$, is below a threshold $\epsilon$, and 0 otherwise.
3.3 Affine-Invariant Camera Pose
The model’s permutation equivariance, combined with the inherent scale ambiguity of multi-view reconstruction, implies that the output camera poses $(\hat{\mathbf{T}}_{1},\dots,\hat{\mathbf{T}}_{N})$ are only defined up to an arbitrary similarity transformation. This specific type of affine transformation consists of a rigid transformation and a single, unknown global scale factor.
To resolve the ambiguity of the global reference frame, we supervise the network on the relative poses between views. The predicted relative pose $\hat{\mathbf{T}}_{i\leftarrow j}$ from view $j$ to $i$ is computed as:
$\hat{\mathbf{T}}_{i\leftarrow j}=\hat{\mathbf{T}}_{i}^{-1}\hat{\mathbf{T}}_{j}$ (7)
Each predicted relative pose $\hat{\mathbf{T}}_{i\leftarrow j}$ is composed of a rotation $\hat{\mathbf{R}}_{i\leftarrow j}\in SO(3)$ and a translation $\hat{\mathbf{t}}_{i\leftarrow j}\in\mathbb{R}^{3}$. While the relative rotation is invariant to this global transformation, the relative translation’s magnitude is ambiguous. We resolve this by leveraging the optimal scale factor, $s^{*}$, that is computed by aligning the predicted point map to the ground truth (as detailed in a previous section). This single, consistent scale factor is used to rectify all predicted camera translations, allowing us to directly supervise both the rotation and the correctly-scaled translation components.
The camera loss $\mathcal{L}_{\text{cam}}$ is a weighted sum of a rotation loss term and a translation loss term, averaged over all ordered view pairs where $i\neq j$:
$\mathcal{L}_{\text{cam}}=\frac{1}{N(N-1)}\sum_{i\neq j}\mathopen{}\mathclose{{\left(\mathcal{L}_{\text{rot}}(i,j)+\lambda_{trans}\mathcal{L}_{\text{trans}}(i,j)}}\right)$ (8)
where $\lambda$ is a hyperparameter to balance the two terms.
The rotation loss minimizes the geodesic distance (angle) between the predicted relative rotation $\hat{\mathbf{R}}_{i\leftarrow j}$ and its ground-truth target $\mathbf{R}_{i\leftarrow j}$:
$\mathcal{L}_{\text{rot}}(i,j)=\arccos\mathopen{}\mathclose{{\left(\frac{\text{Tr}\mathopen{}\mathclose{{\left((\mathbf{R}_{i\leftarrow j})^{\top}\hat{\mathbf{R}}_{i\leftarrow j}}}\right)-1}{2}}}\right)$ (9)
For the translation loss, we compare our scaled prediction against the ground-truth relative translation, $\mathbf{t}_{i\leftarrow j}$. We use the Huber loss, $\mathcal{H}_{\delta}$, for its robustness to outliers:
$\mathcal{L}_{\text{trans}}(i,j)=\mathcal{H}_{\delta}(s^{*}\hat{\mathbf{t}}_{i\leftarrow j}-\mathbf{t}_{i\leftarrow j})$ (10)
Our affine-invariant camera model builds on a key insight: real-world camera paths are highly structured, not random. They typically lie on a low-dimensional manifold—for instance, a camera orbiting an object moves along a sphere, while a car-mounted camera follows a curve.
We quantitatively analyze the structure of the predicted pose distributions in Fig. 4. The eigenvalue analysis confirms that the variance of our predicted poses is concentrated along significantly fewer principal components than VGGT, validating the low-dimensional structure of our output. We discuss this further in Appendix A.3.
Figure 4: Comparison of predicted pose distributions. Our predicted pose distribution exhibits a clear low-dimensional structure.
3.4 Model Training
Our model is trained end-to-end by minimizing a composite loss function, $\mathcal{L}$, which is a weighted sum of the point reconstruction loss, the confidence loss, and the camera pose loss:
$\mathcal{L}=\mathcal{L}_{\text{points}}+\lambda_{\text{normal}}\mathcal{L}_{\text{normal}}+\lambda_{\text{conf}}\mathcal{L}_{\text{conf}}+\lambda_{\text{cam}}\mathcal{L}_{\text{cam}}$ (11)
To ensure robustness and wide applicability, we train the model on a large-scale aggregation of 15 diverse datasets. This combined dataset provides extensive coverage of both indoor and outdoor environments, encompassing a wide variety of scenes from synthetic renderings to real-world captures. The specific datasets include GTA-SfM [35], CO3D [21], WildRGB-D [41], Habitat [23], ARKitScenes [2], TartanAir [40], ScanNet [5], ScanNet++ [44], BlendedMVG [43], MatrixCity [15], MegaDepth [16], Hypersim [22], Taskonomy [45], Mid-Air [8], and an internal dynamic scene dataset. Details of model training can be found in Appendix A.2.
4 Experiments
We evaluate our method on four tasks: camera pose estimation (Sec. 4.1), point map estimation (Sec. 4.2), video depth estimation (Sec. 4.3) and monocular depth estimation (Sec. 4.4). Across all tasks, our method achieves state-of-the-art(SOTA) or comparable performance against existing feed-forward 3D reconstruction methods. To validate the effectiveness of our design, We also conduct several analyses: (1) a robustness evaluation against input image sequence permutations (Sec. 4.5), (2) an ablation study on scale-invariant point maps and affine-invariant camera poses (Sec. 4.6).
4.1 Camera Pose Estimation
We assess predicted camera pose using two distinct sets of metrics: angular accuracy (following [33, 39, 34]) and distance error (following [48, 46, 36]).
Angular Accuracy Metrics. Following prior work [39, 34], we evaluate predicted camera poses on the scene-level RealEstate10K [49] and object-centric Co3Dv2 [21] datasets, both featuring over 1000 test sequences. For each sequence, we randomly sample 10 images, form all possible pairs, and compute the angular errors of the relative rotation and translation vectors. This process yields the Relative Rotation Accuracy (RRA) and Relative Translation Accuracy (RTA) at a given threshold (e.g., RRA@30 for 30 degrees). The Area Under the Curve (AUC) of the min(RRA,RTA)-threshold curve serves as a unified metric. As shown in Tab. 1, our method sets a new SOTA benchmark in zero-shot generalization on RealEstate10K and achieves performance comparable to the SOTA on the in-domain Co3Dv2 dataset. These results underscore our model’s strong generalization capabilities while maintaining excellent performance on familiar data distributions.
Distance Error Metrics. Following [36], we report the Absolute Trajectory Error (ATE), Relative Pose Error for translation (RPE trans), and Relative Pose Error for rotation (RPE rot) on the synthetic outdoor Sintel [3] dataset, as well as the real-world indoor TUM-dynamics [29] and ScanNet [5] datasets. Predicted camera trajectories are aligned with the ground truth via a Sim(3) transformation before calculating the errors. The results in Tab. 2 show that our method significantly outperforms other approaches on Sintel while achieving competitive SOTA results alongside VGGT on TUM-Dynamics and ScanNet.
Figure 5: Qualitative comparison of multi-view 3D reconstruction. Compared to other multi-frame feed-forward reconstruction methods, $\pi^{3}$ produces cleaner, more accurate and more complete reconstructions with fewer artifacts.
Table 1: Camera Pose Estimation on RealEstate10K [49] and Co3Dv2 [21]. Metrics measure the ratio of angular accuracy of rotation/translation under an error of 30 degrees, the higher the better. All methods have seen Co3Dv2 samples during training time, while RealEstate10K is excluded from trainset except for CUT3R.
Method RealEstate10K (unseen) Co3Dv2
RRA@30 $\uparrow$ RTA@30 $\uparrow$ AUC@30 $\uparrow$ RRA@30 $\uparrow$ RTA@30 $\uparrow$ AUC@30 $\uparrow$
Fast3R [42] 99.05 81.86 61.68 97.49 91.11 73.43
CUT3R [36] 99.82 95.10 81.47 96.19 92.69 75.82
FLARE [47] 99.69 95.23 80.01 96.38 93.76 73.99
VGGT [34] 99.97 93.13 77.62 98.96 97.13 88.59
$\pi^{3}$ (Ours) 99.99 95.62 85.90 99.05 97.33 88.41
Table 2: Camera Pose Estimation on Sintel [3], TUM-dynamics [29] and ScanNet [5]. Metrics measure the distance error of rotation/translation, the lower the better. All methods except Aether have seen ScanNet or ScanNet++ [44] samples during training time. Zero-shot pose estimation accuracy is evaluated on Sintel and TUM-dynamics for all methods.
Method Sintel TUM-dynamics ScanNet (seen)
[…]
Fast3R [42] 0.371 0.298 13.75 0.090 0.101 1.425 0.155 0.123 3.491
CUT3R [36] 0.217 0.070 0.636 0.047 0.015 0.451 0.094 0.022 0.629
Aether [31] 0.189 0.054 0.694 0.092 0.012 1.106 0.176 0.028 1.204
FLARE [47] 0.207 0.090 3.015 0.026 0.013 0.475 0.064 0.023 0.971
VGGT [34] 0.167 0.062 0.491 0.012 0.010 0.311 0.035 0.015 0.382
$\pi^{3}$ (Ours) 0.074 0.040 0.282 0.014 0.009 0.312 0.031 0.013 0.347
Table 3: Point Map Estimation on DTU [12] and ETH3D [26]. Keyframes are selected every 5 images.
Method DTU ETH3D
Acc. $\downarrow$ Comp. $\downarrow$ N.C. $\uparrow$ Acc. $\downarrow$ Comp. $\downarrow$ N.C. $\uparrow$
Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med.
Fast3R [42] 3.340 1.919 2.929 1.125 0.671 0.755 0.832 0.691 0.978 0.683 0.667 0.766
CUT3R [36] 4.742 2.600 3.400 1.316 0.679 0.764 0.617 0.525 0.747 0.579 0.754 0.848
FLARE [47] 2.541 1.468 3.174 1.420 0.684 0.774 0.464 0.338 0.664 0.395 0.744 0.864
VGGT [34] 1.338 0.779 1.896 0.992 0.676 0.766 0.280 0.185 0.305 0.182 0.853 0.950
$\pi^{3}$ (Ours) 1.198 0.646 1.849 0.607 0.678 0.768 0.194 0.131 0.210 0.128 0.883 0.969
Table 4: Point Map Estimation on 7-Scenes [27] and NRGBD [1]. Keyframes are selected every 200 images (for 7-Scenes) and 500 images (for NRGBD) for sparse view, and every 40 images (for 7-Scenes) and 100 images (for NRGBD) for dense view.
Method View 7-Scenes NRGBD
Acc. $\downarrow$ Comp. $\downarrow$ NC. $\uparrow$ Acc. $\downarrow$ Comp. $\downarrow$ NC. $\uparrow$
Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med.
Fast3R [42] sparse 0.095 0.065 0.144 0.089 0.673 0.759 0.135 0.091 0.163 0.104 0.759 0.877
CUT3R [36] 0.093 0.049 0.102 0.051 0.704 0.805 0.104 0.041 0.079 0.031 0.822 0.968
FLARE [47] 0.085 0.057 0.145 0.107 0.696 0.780 0.053 0.024 0.051 0.025 0.877 0.988
VGGT [34] 0.044 0.025 0.056 0.033 0.733 0.845 0.051 0.029 0.066 0.038 0.890 0.981
$\pi^{3}$ (Ours) 0.047 0.029 0.075 0.049 0.742 0.841 0.026 0.015 0.028 0.014 0.916 0.992
Fast3R [42] dense 0.040 0.017 0.056 0.018 0.644 0.725 0.072 0.030 0.050 0.016 0.790 0.934
CUT3R [36] 0.023 0.010 0.027 0.008 0.669 0.764 0.086 0.037 0.048 0.017 0.800 0.953
FLARE [47] 0.019 0.007 0.026 0.013 0.684 0.785 0.023 0.011 0.018 0.008 0.882 0.986
VGGT [34] 0.022 0.008 0.026 0.012 0.666 0.760 0.017 0.010 0.015 0.005 0.893 0.988
$\pi^{3}$ (Ours) 0.016 0.007 0.022 0.011 0.689 0.792 0.015 0.008 0.013 0.005 0.898 0.987
4.2 Point Map Estimation
Following the evaluation settings in [36], we evaluate the quality of reconstructed multi-view point maps on the scene-level 7-Scenes [27] and NRGBD [1] datasets under both sparse and dense view conditions. For sparse views, keyframes are sampled with a stride of 200 (7-Scenes) or 500 (NRGBD), while for dense views, the stride is reduced to 40 (7-Scenes) or 100 (NRGBD). We also extend our evaluation to the object-centric DTU [12] and scene-level ETH3D [26] datasets, sampling keyframes every 5 images. Predicted point maps are aligned to the ground truth using the Umeyama algorithm for a coarse Sim(3) alignment, followed by refinement with the Iterative Closest Point (ICP) algorithm.
Consistent with prior works [1, 39, 32, 36], we report Accuracy (Acc.), Completion (Comp.), and Normal Consistency (N.C.) in Tab. 3 and Tab. 4. These results highlight the effectiveness of our method in a broad spectrum of 3D reconstruction tasks, spanning object-level and scene-level cases (Tab. 3), and demonstrating robustness on both real-world and synthetic datasets (Tab. 4).
To provide a comprehensive evaluation, we further analyze performance under sparse-view and dense-view conditions using the 7-Scenes and NRGBD datasets (Tab. 4). The sparse-view setup, defined by limited inter-frame overlap, presents a highly ill-posed problem requiring the model to exploit strong spatial priors. For completeness, we also consider the dense-view scenario, where ample observations facilitate reconstruction. The results confirm that our method achieves consistently robust performance in both challenging sparse-view and favorable dense-view settings.
4.3 Video Depth Estimation
Following the methodology of [46, 36], we evaluate our method on the task of video depth estimation using the Sintel [3], Bonn [19], and KITTI [10] datasets. We report the Absolute Relative Error (Abs Rel) and the prediction accuracy at a threshold of $\delta<1.25$. The metrics are evaluated under two alignment settings: (i) scale-only alignment and (ii) joint scale and 3D translation alignment.
As reported in Tab. 5, our method achieves a new state-of-the-art performance across all three datasets and both alignment settings within feed-forward 3D reconstruction methods. Notably, it also delivers exceptional efficiency, running at 57.4 FPS on KITTI, significantly faster than VGGT (43.2 FPS) and Aether (6.14 FPS), despite having a smaller model size.
Table 5: Video Depth Estimation on Sintel [3], Bonn [19] and KITTI [10]. FPS is evaluated on KITTI using one A800 GPU.
Method Params Align Sintel Bonn KITTI FPS
Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$
DUSt3R [39] 571M scale 0.662 0.434 0.151 0.839 0.143 0.814 1.25
MASt3R [13] 689M 0.558 0.487 0.188 0.765 0.115 0.848 1.01
MonST3R [46] 571M 0.399 0.519 0.072 0.957 0.107 0.884 1.27
Fast3R [42] 648M 0.638 0.422 0.194 0.772 0.138 0.834 65.8
MVDUSt3R [30] 661M 0.805 0.283 0.426 0.357 0.456 0.342 0.69
CUT3R [36] 793M 0.417 0.507 0.078 0.937 0.122 0.876 6.98
Aether [31] 5.57B 0.324 0.502 0.273 0.594 0.056 0.978 6.14
FLARE [47] 1.40B 0.729 0.336 0.152 0.790 0.356 0.570 1.75
VGGT [34] 1.26B 0.299 0.638 0.057 0.966 0.062 0.969 43.2
$\pi^{3}$ (Ours) 959M 0.233 0.664 0.049 0.975 0.038 0.986 57.4
DUSt3R [39] 571M scale & shift 0.570 0.493 0.152 0.835 0.135 0.818 1.25
MASt3R [13] 689M 0.480 0.517 0.189 0.771 0.115 0.849 1.01
MonST3R [46] 571M 0.402 0.526 0.070 0.958 0.098 0.883 1.27
Fast3R [42] 648M 0.518 0.486 0.196 0.768 0.139 0.808 65.8
MVDUSt3R [30] 661M 0.619 0.332 0.482 0.357 0.401 0.355 0.69
CUT3R [36] 793M 0.534 0.558 0.075 0.943 0.111 0.883 6.98
Aether [31] 5.57B 0.314 0.604 0.308 0.602 0.054 0.977 6.14
FLARE [47] 1.40B 0.791 0.358 0.142 0.797 0.357 0.579 1.75
VGGT [34] 1.26B 0.230 0.678 0.052 0.969 0.052 0.968 43.2
$\pi^{3}$ (Ours) 959M 0.210 0.726 0.043 0.975 0.037 0.985 57.4
Figure 6: Qualitative comparison of in-the-wild multi-view 3D reconstruction. $\pi^{3}$ demonstrates superior robustness on challenging in-the-wild sequences, consistently producing more coherent and complete 3D structures for both dynamic and complex static scenes compared to other feed-forward approaches.
4.4 Monocular Depth Estimation
Similar to our video depth evaluation, we compare our predicted monocular depth against other feed-forward reconstruction methods. Following [46, 36], we use four datasets to evaluate the accuracy of scale-invariant monocular depth. We continue to use the Absolute Relative Error (Abs Rel) and threshold accuracy ($\delta$<1.25) as metrics. However, in this setting, each depth map is aligned independently with its ground truth, in contrast to the video depth evaluation, where a single scale (and shift) factor is applied to the entire image sequence.
As reported in Tab. 6, our method achieves state-of-the-art results among multi-frame feed-forward reconstruction approaches, even though it is not explicitly optimized for single-frame depth estimation. Notably, it performs competitively with MoGe [37, 38], one of the top-performing monocular depth estimation models.
Table 6: Monocular Depth Estimation on Sintel [3], Bonn [19], KITTI [10] and NYU-v2 [28].
Method Sintel Bonn KITTI NYU-v2
[…]
DUSt3R [39] 0.488 0.532 0.139 0.832 0.109 0.873 0.081 0.909
MASt3R [13] 0.413 0.569 0.123 0.833 0.077 0.948 0.110 0.865
MonST3R [46] 0.402 0.525 0.069 0.954 0.098 0.895 0.094 0.887
Fast3R [42] 0.544 0.509 0.169 0.796 0.120 0.861 0.093 0.898
CUT3R [36] 0.418 0.520 0.058 0.967 0.097 0.914 0.081 0.914
FLARE [47] 0.606 0.402 0.130 0.836 0.312 0.513 0.089 0.898
VGGT [34] 0.335 0.599 0.053 0.970 0.082 0.947 0.056 0.951
MoGe 0.273 0.695 0.050 0.976 0.049 0.979 0.055 0.952
- v1 [37]  - 0.273  - 0.695  - 0.050  - 0.976  - 0.054  - 0.977  - 0.055  - 0.952
- v2 [38]  - 0.277  - 0.687  - 0.063  - 0.973  - 0.049  - 0.979  - 0.060  - 0.940
$\pi^{3}$ (Ours) 0.277 0.614 0.044 0.976 0.060 0.971 0.054 0.956
4.5 Robustness Evaluation
A key property of our proposed architecture is permutation equivariance, ensuring that its outputs are robust to variations in the input image sequence order. To empirically verify this, we conduct experiments on the DTU [12] and ETH3D [26] datasets. For each input sequence of length $N$, we perform $N$-fold separate inferences, where in each run we replace the original first frame with a different frame from the sequence. We then compute the standard deviation of the reconstruction metrics across these $N$ outputs. A lower standard deviation indicates higher robustness to input order variations.
As reported in Tab. 7, our method achieves near-zero standard deviation across all metrics on DTU and ETH3D, outperforming existing approaches by several orders of magnitude. For instance, on DTU, our mean accuracy standard deviation is 0.003, while VGGT reports 0.033. On ETH3D, our model achieves effectively zero variance. This stark contrast highlights the limitations of reference-frame-dependent methods, which exhibit significant sensitivity to input order. Our results provide compelling evidence that the proposed architecture is genuinely permutation-equivariant, ensuring consistent and order-independent 3D reconstruction.
Table 7: Standard Deviation of Point Cloud Estimation on DTU [12] and ETH3D [26].
Method DTU ETH3D
Acc. std. $\downarrow$ Comp. std. $\downarrow$ N.C. std. $\downarrow$ Acc. std. $\downarrow$ Comp. std. $\downarrow$ N.C. std. $\downarrow$
Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med.
Fast3R [42] 0.578 0.451 0.677 0.376 0.007 0.009 0.182 0.205 0.381 0.273 0.047 0.072
CUT3R [36] 1.750 1.047 1.748 1.273 0.013 0.017 0.214 0.225 0.430 0.391 0.062 0.074
FLARE [47] 0.720 0.494 1.346 1.134 0.009 0.012 0.171 0.187 0.251 0.188 0.048 0.053
VGGT [34] 0.033 0.022 0.054 0.036 0.007 0.007 0.049 0.040 0.062 0.042 0.022 0.015
$\pi^{3}$ (Ours) 0.003 0.002 0.006 0.003 0.001 0.001 0.000 0.000 0.000 0.000 0.001 0.000
4.6 Ablation Study
Table 8: Ablation study on the key components of our model. We show how the performance metric improves as each component is added to the baseline.
Model ETH3D 7-Scenes NRGBD
Acc. $\downarrow$ Comp. $\downarrow$ N.C $\uparrow$ Acc. $\downarrow$ Comp. $\downarrow$ N.C $\uparrow$ Acc. $\downarrow$ Comp. $\downarrow$ N.C $\uparrow$
Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med. Mean Med.
Model 1 0.229 0.150 0.166 0.103 0.802 0.930 0.020 0.010 0.019 0.009 0.715 0.834 0.034 0.018 0.025 0.011 0.859 0.977
Model 2 0.197 0.118 0.118 0.065 0.820 0.943 0.020 0.009 0.020 0.008 0.716 0.837 0.031 0.018 0.023 0.010 0.861 0.978
Full Model 0.131 0.076 0.079 0.043 0.841 0.957 0.019 0.009 0.020 0.009 0.723 0.843 0.028 0.015 0.022 0.010 0.875 0.981
To validate the effectiveness of our proposed components, we conducted an ablation study by systematically removing features from our complete model. First, we created Model 2 by removing the affine-invariant camera pose modeling from our full model. Subsequently, we derived Model 1 by also removing the scale-invariant pointmap modeling from Model 2.
The primary difference between our full model and the ablated models (Model 1 and Model 2) is that the latter two incorporate a camera token. This token is essential for distinguishing the reference view, as the model is no longer permutation-equivariant after the removal of the affine-invariant camera pose modeling. At each iteration, the camera token is concatenated with a randomly selected reference view before the alternating-attention module similar to [34]. We compute an angle loss for rotation and a Huber loss for translation between the predicted and ground-truth poses in the reference view’s coordinate system for Model 1 and Model 2. While Model 1 and Model 2 share an identical architecture and parameter count, their key distinctions lie in the loss calculation and normalization processes. For Model 1, we neither perform alignment during the loss computation for the predicted pointmap nor do we normalize the pointmap itself. We found that applying normalization in this specific case led to anomalous and significantly degraded performance, a phenomenon also observed in prior work [34]. In contrast, the predicted local pointmaps are normalized for both Model 2 and the full model.
For a fair comparison, all models were trained for 80 epochs, with 800 iterations per epoch, on images with a resolution of $224\times 224$. They shared the same initialization procedure as our final model: we loaded pre-trained weights for the VGGT encoder and alternating-attention layers, and kept the encoder frozen throughout training. The comparative results for pointmap estimation across three datasets are presented in Table 8. For the 7-Scenes and NRGBD datasets, we use the same dense view setting as in the previous section.
We found that scale-invariant pointmap modeling does not yield significant performance gains on indoor datasets like 7-Scenes and NRGBD. For outdoor data, however, the performance improvement is substantially more pronounced. This observation is consistent with previous studies on scale-invariant depth, which have shown that outdoor scenes are more significantly affected by scale ambiguity. Furthermore, we observed that affine-invariant camera pose modeling consistently enhances the final performance. More importantly, unlike Model 1 and Model 2, its inclusion renders the model permutation-equivariant. Consequently, the model becomes robust to both the order of input frames and the selection of the reference view.
5 Conclusion
In this work, we introduced $\pi^{3}$, a feed-forward neural network that presents a new paradigm for visual geometry reconstruction by eliminating the reliance on a fixed reference view. By leveraging a fully permutation-equivariant architecture, our model is inherently robust to input ordering and leads to higher accuracy. This design choice removes a critical inductive bias found in previous methods, allowing our simple yet powerful approach to achieve state-of-the-art performance on a wide array of tasks, including camera pose estimation, depth estimation, and dense reconstruction. $\pi^{3}$ demonstrates that reference-free systems are not only viable but can lead to more stable and versatile 3D vision models.
Appendix A Appendix
A.1 Architecture Details
The encoder and alternating attention modules are the same as those in VGGT [34], with the exception that we use only 36 layers for the alternating attention module, whereas VGGT uses 48. The decoders for camera poses, local point maps, and confidence scores share the same architecture but do not share weights. This architecture is a lightweight, 5-layer transformer that applies self-attention exclusively to the features of each individual image. Following the decoder, the output heads vary by task. The heads for local point maps and confidence scores consist of a simple MLP followed by a pixel shuffle operation. For camera poses, the head is adapted from Reloc3r [6] and uses an MLP, average pooling, and another MLP. The rotation is initially predicted in a 9D representation [14] and is then converted to a 3×3 rotation matrix via SVD orthogonalization.
A.2 Training Details
We train $\pi^{3}$ in two stages, a process similar to Dust3R [39]. First, the model is trained on a low resolution of 224 $\times$ 224 pixels. Then, it is fine-tuned on images of random resolutions where the total pixel count is between 100,000 and 255,000 and the aspect ratio is sampled from the range [0.5, 2.0], a strategy similar to MoGe [37]. We use a dynamic batch sizing strategy similar to VGGT. In the first stage, we sample 64 images per GPU, and in the second stage, we sample 48 images per GPU. Each batch is composed of 2 to 24 images. Each training stage runs for 80 epochs, with each epoch comprising 800 iterations. Our final model is not trained from scratch. Instead, we initialize the weights for the encoder and the alternating attention module from the pre-trained VGGT model, and we keep the encoder frozen during training. We train the first stage on 16 A100 GPUs and the second stage on 64 A100 GPUs. For our loss function, we set the weights for each component as follows: $\lambda_{\text{normal}}=1.0$, $\lambda_{\text{conf}}=0.05$, $\lambda_{\text{cam}}=0.1$, and $\lambda_{\text{trans}}=100.0$. The implementation of our normal loss follows that of MoGe, and the resolution for aligning the local point map loss is set to 4096. Regarding optimization, we assign different initial learning rates to model components: $5\times 10^{-6}$ for the encoder and $5\times 10^{-5}$ for all other modules. We employ a OneCycleLR scheduler, where the learning rate anneals from its maximum value down to a minimal value over the entire training duration following a cosine curve. We use the same learning rate and scheduler settings for both stages. The confidence head is not trained jointly with the other modules. Instead, after completing the two main training stages, we freeze the rest of the network and train the confidence head in isolation. This final stage converges rapidly, typically within a few epochs, without impacting the model’s overall performance. We use gradient clipping with a norm of 1.0.
Figure 7: Comparison of predicted pose distributions. We visualize the predicted pose distributions in 3D space. $\pi^{3}$ shows a clear low-dimensional structure, while VGGT’s distribution is scattered.
A.3 Discussion for Predicted Pose Distribution
In Fig. 7, we present a 3D visualization of the camera pose distributions predicted by our method and by VGGT. To visualize the rotational component, we map it to the RGB color space. It is clear that our predictions form a distinct low-dimensional structure, while the distribution from VGGT is much more scattered and random. That may be one of the potential reasons for the fast speed of training convergence.
A.4 Comparison with VGGT
This section details an experiment designed solely for a fair comparison against VGGT [34]. A direct comparison is challenging because training our model from scratch with only its core objectives (camera poses and local pointmaps) leads to suboptimal convergence, whereas VGGT’s design incorporates a multi-task learning setup.
For a fair comparison, and similar to VGGT, we also predict a global pointmap to serve as a regularizer. We temporarily adapt our training to mirror VGGT’s methodology. We introduce an auxiliary head to predict a global pointmap relative to a reference frame, using a loss analogous to Eq. 5. We directly use the scale factor from the alignment of local pointmaps. Please note that the reference view is incorporated as context via cross-attention, exclusively within the global pointmap head. This head only serves as a regularization term and our final model remains fully permutation-equivariant.
We train both our adapted model and VGGT under these identical, multi-task conditions: from scratch (except for DINOv2 encoders) on the same data, at a $224\times 224$ resolution for 80 epochs (800 steps/epoch). We use the same data as described in Section 3.4.
Table 9: Comparison with VGGT when trained from scratch.
Method ETH3D 7-Scenes NRGB
Acc. $\downarrow$ Comp. $\downarrow$ Acc. $\downarrow$ Comp. $\downarrow$ Acc. $\downarrow$ Comp. $\downarrow$
VGGT 0.563 0.449 0.057 0.046 0.060 0.042
$\pi^{3}$ (ours) 0.418 0.266 0.059 0.071 0.052 0.035
As shown in Table 9, $\pi^{3}$ outperforms VGGT on two of the three benchmarks. It is important to mention that for this VGGT baseline, we did not utilize its tracking branch, as the official implementation did not provide clear instructions or clean code for its usage.
A.5 Limitations
Our model demonstrates strong performance, but it also has several key limitations. First, it is unable to handle transparent objects, as our model does not explicitly account for complex light transport phenomena. Second, compared to contemporary diffusion-based approaches, our reconstructed geometry lacks the same level of fine-grained detail. Finally, the point cloud generation relies on a simple upsampling mechanism using an MLP with pixel shuffling. While efficient, this design can introduce noticeable grid-like artifacts, particularly in regions with high reconstruction uncertainty.